A 620-gram salt-and-water solution contains 50 g of salt. This mixture is left in the open air, and 120 g of water evaporates from the solution. What is the percent concentration of salt in the remaining solution?
Answer by stanbon(75887)

A 620-gram salt-and-water solution contains 50 g of salt. This mixture is left in the open air, and 120 g of water evaporates from the solution. What is the percent concentration of salt in the remaining solution?
---------
Equation: water - water = water
(620-50) - 120 = 450 g of water
------
Ans: 50/450 = 1/9 = 11.1% salt
===============
Cheers,
Stan H.
===============